# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v1

@@ Appendix D ContextFlow Algorithm
19:   Obtain spatial-prior-aware OT couplings $\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\leftarrow\operatorname{diag}(\bm{u})\mathbf{K}\operatorname{diag}(\bm{v})$
20:   Sample $t\sim\mathcal{U}(t_{i},t_{i+1})$ and $\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\}$
21:   Sample $\bm{x}_{t}\sim\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right)$
22:   $L_{\mathrm{CFM}}\leftarrow\frac{1}{|\mathcal{B}|}\sum_{t,(\bm{x}_{i},\bm{x}_{i+1})}\left\|u_{\theta}(\bm{x}_{t},t)-\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}}\right\|_{2}^{2}$
23:  end for
24:  $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}L_{\mathrm{CFM}}$
25:end while
Appendix E Kantarovich-OT Formulation
Appendix F Evaluation Metrics
F.1 2-Wasserstein
The 2-Wasserstein distance ($\mathcal{W}_{2}$ between empirical distributions $\mu,\nu$ is defined as:
where $\Pi(\mu,\nu)$ denotes the set of couplings between $\mu$ and $\nu$.
F.2 Weighted 2-Wasserstein
Implausible velocity fields can steer a cell’s transcriptional trajectory in unrealistic directions, potentially leading to entirely different terminal cell types. We thus employ the weighted 2-Wasserstein metric, which ensures the evaluation accounts for both transcriptional similarity and the distributional balance of cell types. We define the weighted 2-Wasserstein distance (Weighted $\mathcal{W}_{2}$) between true and predicted distributions as:
where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples. To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model (Chen & Guestrin, 2016) trained for each dataset.
F.3 Energy Distance
Let $\mu$ and $\nu$ be probability distributions with samples $X=\{\bm{x}_{i}\}_{i=1}^{m}\sim\mu$ and $Y=\{\bm{y}_{j}\}_{j=1}^{n}\sim\nu$. The squared empirical energy distance (Energy) is defined as:
F.4 Maximum Mean Discrepancy
For the same samples, the unbiased empirical estimate of the squared maximum mean discrepancy (MMD) with kernel $\kappa$ is defined as:
Appendix G Biological Priors and Dataset Visualizations
G.1 (Im-)Plausibility of OT-Couplings
To demonstrate the need for integrating biological priors within a generative framework, we computed the Entropic-OT plan (Section 2.3) for the MOTFM framework and the PAER-OT plan (Section 3.3) for the ContextFlow framework. From these transport plans, we sampled couplings corresponding to the first two stages of the Brain Regeneration dataset (Wei et al., 2022) together with their associated cell types. Figures 3(a) and 3(b) illustrate the Excitatory–Inhibitory lineage switches present in these sampled couplings. Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
In our transport plan couplings, we observed the following cell type lineage switches:
  • Immature MSN $\rightarrow$ Immature nptxEX
  • Immature MSN $\rightarrow$ Immature dpEX
  • Immature MSN $\rightarrow$ Immature CMPN
  • Immature nptxEX $\rightarrow$ Immature cckIN
  • Immature nptxEX $\rightarrow$ Immature MSN
@@ Appendix G Biological Priors and Dataset Visualizations
Of these, $54$ implausible transitions arose from the Entropic-OT plan compared to the $24$ under the PAER-OT plan, with the specific transitions detailed in the figure legends. We also observed that the Entropic-OT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right. In contrast, the PAER-OT formulation typically restricted transitions to within the same hemisphere, reflecting its integration of spatially aware contextual information. These observations provide strong motivation for incorporating biological priors through ContextFlow as a principled approach to learning biologically consistent developmental trajectories.
@@ G.2 Cell type distributions over time
Figures 46 present the spatial maps of the transcriptomics datasets across different time points, illustrating how tissue organization and cell type distributions evolve during development and regeneration. These maps highlight not only changes in cellular composition but also the preservation of spatial neighborhoods and geometrical arrangements of specific cell types over time. Such contextual information, specific to spatial transcriptomics, remains inaccessible to standard flow-matching frameworks. By contrast, ContextFlow is designed to exploit these spatial features, enabling the inference of trajectories that are both temporally smooth and spatially coherent.
@@ G.3 Ligand Receptor Interactions
Based on the activation of NPTX2–NPTXR in Figure 7, we observe that the corresponding communication pattern naturally biases the optimal couplings towards transitions such as Immature dpEX $\rightarrow$ dpEX and Immature nptxEX $\rightarrow$ nptxEX (Figure 8). These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context. This example highlights the richness of the contextual information captured by our proposed biological prior, and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process leads to more interpretable and biologically consistent trajectories.
Figure 8: Visual translation of the bias that NPTX2–NPTXR LR pattern provides in terms of cell type coupling for the two consecutive slides.

### v2

@@ E.1 Algorithm Pseudocode
20:   Obtain spatial-prior-aware OT couplings $\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\leftarrow\operatorname{diag}(\bm{u})\mathbf{K}\operatorname{diag}(\bm{v})$
21:   Sample $t\sim\mathcal{U}(t_{i},t_{i+1})$ and $\{(\bm{x}_{i},\bm{x}_{i+1}):(\bm{x}_{i},\bm{x}_{i+1})\sim\mathbf{\Pi}^{\mathrm{CTF}}_{i,i+1}\}$
22:   Sample $\bm{x}_{t}\sim\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right)$
23:   $L_{\mathrm{CFM}}\leftarrow\frac{1}{|\mathcal{B}|}\sum_{t,(\bm{x}_{i},\bm{x}_{i+1})}\left\|u_{\theta}(\bm{x}_{t},t)-\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}}\right\|_{2}^{2}$
24:  end for
25:  $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}L_{\mathrm{CFM}}$
26:end while
E.2 Time Complexity Analysis
The training time of ContextFlow is comparable to that of Minibatch-OT FM (Tong et al., 2024) (Figure 3), as both methods solve an entropic variant of optimal transport using the GPU-optimized Sinkhorn algorithm, alongside forward and backward propagation steps that are also GPU-accelerated. Although ContextFlow incorporates prior knowledge, such as spatial smoothness (Equation 9) and cell-cell communication patterns (Equation 10), their corresponding features are computed once during preprocessing, resulting in a one-time cost. The precomputed features can be reused across multiple hyperparameter settings and model variants, making ContextFlow highly scalable and efficient.
Data Preprocessing. The following preprocessing steps generate additional biologically informed features that complement the original transcriptomic profiles. These features incur a one-time computational cost and can be reused across different experiments and model configurations.
Spatial Smoothness (SS). We employ a nearest neighbor (NN) algorithm for calculating the mean of local transcriptomic features for each cell. The computational complexity of the NN search is known to be $O(N^{2}d)$, where $N$ denotes the total number of points considered and $d$ denotes the data dimension.
Cell-Cell Communication Patterns (LR). We employ spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), for computing LR features, where we applied the cosine similarity metric to gene expression profiles and used the recommended hyperparameters. The exact runtime complexity for LIANA+ is unknown. Table 6 summarizes the total time taken for the Brain Regeneration and Mouse Organogenesis datasets in our case.
Table 6: Runtime for computing cell-cell communication patterns.
Dataset Total Number of Cells Runtime (seconds)
Brain Regeneration (Wei et al., 2022) 28,780 23.35
Mouse Organogenesis (Chen et al., 2022) 399,248 200.40
Training of ContextFlow. The training time complexity largely depends on the total number of training epochs ($E$), mini-batch size ($B$), time per forward and backward pass ($P$), transcriptomic feature dimension ($d$), and total LR pairs ($l$). Below, we compute the time complexity for each step in ContextFlow:
TPM Construction. The construction of the transition plausibility matrix involves a calculation of pairwise distances for each mini-batch, resulting in the runtime of $O(B^{2}(l+d))$.
Sinkhorn Algorithm. According to Theorem 3.1, we know that Sinkhorn iterations can be adapted to solve the prior-aware entropy regularization problem (Equation 13). Since the Sinkhorn algorithm has a well-known quadratic time complexity (Cuturi, 2013), the runtime for computing minibatch OT couplings in ContextFlow is $O(B^{2})$.
Figure 3: Training time comparisons for different algorithms with a mini-batch size of $256$ on the Brain Regeneration dataset.
Total Runtime. Putting pieces together, across all the training epochs, the total runtime complexity of ContextFlow turns out to be $O(E\times(B^{2}(d+l)+P))$. As shown in Figure 3, the runtime is linearly dependent on the total epochs $E$, with different linear rates for various configurations. CFM is the fastest because it bypasses the optimal transport coupling step required by the other methods.
Appendix F Spatiotemporal Optimal Transport
In this section, we compare the recent state-of-the-art spatiotemporal alignment methods, including DeST-OT (Halmos et al., 2025) and TOAST (Ceccarelli et al., 2025), with our prior-aware entropy regularized (PAER) OT objective used in ContextFlow (CTF-H). It is important to note that these OT methods are not generative models and are only used for pairwise alignment tasks. ContextFlow, on the other hand, is a generative model that learns a dynamic flow across the time horizon and utilizes OT couplings to design better conditional paths for regression. In particular, we compute metrics described in DeST-OT on the Axolotl Brain Regeneration dataset following the same setup as used in flow matching. Specifically, for each time step, we randomly sample a batch of $1000$ cells and compute the corresponding coupling matrix $\mathbf{\Pi}$, which is then used to derive the metrics. We use the CTF-H ($\lambda=0.8$) version of ContextFlow for comparison.
F.1 Metric Comparison
DeST-OT introduces an OT objective for aligning spatial transcriptomic tissue slices from different developmental timesteps, with an emphasis on modeling cell growth and tissue expansion/contraction. The growth distortion metric is designed to assess whether the inferred growth pattern aligns with the changes in cell-type abundance across timesteps. As shown in Table 7, for the growth distortion metric, we find that our CTF-H OT is competitive with DeST-OT and TOAST, despite DeST-OT being specifically developed with consideration for cell growth.
Table 7: Comparison on growth distortion.
Static Pair DeST-OT TOAST CTF-H OT
$1/2$ $0.0007$ $0.0000$ $0.0000$
$2/3$ $0.0042$ $0.0000$ $0.0000$
$3/4$ $0.0027$ $0.0000$ $0.0000$
$4/5$ $0.0009$ $0.0000$ $0.0000$
The migration metric is another important metric introduced in DeST-OT, which measures whether the coupling results in realistic cell movements between timesteps. As seen in Table 8, DeST-OT achieves the best performance, highlighting the advantage of its growth-aware objective compared to TOAST and CTF-H OT, which do not explicitly model tissue expansion or contraction.
Table 8: Comparison on migration.
Static Pair DeST-OT TOAST CTF-H OT
$1/2$ $308.97$ $793.71$ $780.16$
$2/3$ $551.53$ $1103.29$ $1037.29$
$3/4$ $1015.73$ $2052.65$ $1991.43$
$4/5$ $777.13$ $2257.02$ $2222.07$
Lastly, we compute how similar the transcriptomic values of coupled cells are using a Coupled Transcriptomic Distance metric, which is defined as $\sum_{k=1}^{N}\sum_{l=1}^{M}\big\|X_{t_{i}}[k,:]-X_{t_{i+1}}[l,:]\big\|^{2}\times\Pi_{i,j}$, where $X_{t_{i}}[k,:]$ represents the transcriptomic feature of cell $k$ from timestep $t_{i}$ and $X_{t_{i+1}}[l,:]$ represents the transcriptomic feature of cell $l$ from timestep $t_{i+1}$, and $\Pi$ is the OT coupling matrix. From Table 9, we can observe that CTF-H OT is competitive with both DeST and TOAST.
Table 9: On coupled transcriptomic distance.
Static Pair DeST-OT TOAST CTF-H OT
$1/2$ $32.64$ $34.13$ $33.58$
$2/3$ $14.22$ $21.87$ $18.01$
$3/4$ $42.26$ $43.89$ $43.26$
$4/5$ $18.47$ $20.54$ $20.14$
F.2 Runtime Analysis with Varying Sample Size
We also compare the runtime complexity of the above-mentioned OT methods, as shown in Table 10. CTF-H OT is the fastest of the three, followed by DeST-OT and TOAST, and is competitive in the metrics above. We also observe that DeST-OT is the slowest, as expected, since its OT objective includes a Gromov-Wasserstein term, which has $O(n^{3})$ runtime, along with other growth- and tissue-distortion-specific terms.
Table 10: Runtime (s) with varying sample size.
Sample Size DeST-OT TOAST CTF-OT
$10$ $0.0111$ $0.0227$ $0.0009$
$50$ $0.0342$ $0.0114$ $0.0018$
$100$ $0.1892$ $0.0167$ $0.0024$
$150$ $0.4035$ $0.0247$ $0.0053$
$200$ $0.4913$ $0.0337$ $0.0059$
$250$ $0.4571$ $0.0426$ $0.0061$
$300$ $0.6252$ $0.0543$ $0.0074$
$350$ $0.6974$ $0.0656$ $0.0078$
$400$ $0.8612$ $0.0817$ $0.0117$
$450$ $1.1028$ $0.0983$ $0.0110$
$500$ $1.5478$ $0.1197$ $0.0107$
$550$ $1.7448$ $0.1468$ $0.0165$
$600$ $1.9295$ $0.1905$ $0.0209$
$650$ $2.1282$ $0.2077$ $0.0201$
$700$ $2.7013$ $0.2309$ $0.0235$
$750$ $3.2951$ $0.2574$ $0.0327$
$800$ $4.0964$ $0.3001$ $0.0382$
$850$ $4.2001$ $0.3229$ $0.0339$
$900$ $4.4582$ $0.3798$ $0.0483$
$950$ $5.0965$ $0.4206$ $0.0509$
$1000$ $6.1452$ $0.4931$ $0.0375$
Tables 7-10 demonstrate that the design choices of ContextFlow enable it to be highly scalable compared to state-of-the-art spatiotemporal alignment methods, while remaining competitive across several spatiotemporal OT alignment metrics.
@@ F.2 Runtime Analysis with Varying Sample Size
Appendix G Biological Priors and Dataset Visualizations
G.1 (Im-)Plausibility of OT-Couplings
To demonstrate the need of integrating biological priors within a generative framework, we computed the Entropic-OT plan (Section 2.3) for the MOTFM framework and the PAER-OT plan (Section 3.3) for the ContextFlow framework. From these transport plans, we sampled couplings corresponding to the first two stages of the Brain Regeneration dataset (Wei et al., 2022) together with their associated cell types. Figures 4(a) and 4(b) illustrate the Excitatory–Inhibitory lineage switches present in these sampled couplings. Since excitatory and inhibitory neurons have mutually exclusive neurotransmitter functions and originate from distinct progenitor populations with different transcription factor profiles, a transition from excitatory to inhibitory identity is considered biologically implausible.
In our transport plan couplings, we observed the following cell type lineage switches: (i) Immature MSN $\rightarrow$ Immature nptxEX, (ii) Immature MSN $\rightarrow$ Immature dpEX, (iii) Immature MSN $\rightarrow$ Immature CMPN, (iv) Immature nptxEX $\rightarrow$ Immature cckIN, and (v) Immature nptxEX $\rightarrow$ Immature MSN
Of these, $54$ implausible transitions arose from the Entropic-OT plan, compared to $24$ from the PAER-OT plan, with the specific transitions detailed in the figure legends. We also observed that the Entropic-OT formulation produced implausible transitions across brain hemispheres, for example, coupling cells from the left hemisphere with those from the right. In contrast, the PAER-OT formulation typically restricted transitions to within the same hemisphere, reflecting its integration of spatially aware contextual information. These observations provide strong motivation for incorporating biological priors through ContextFlow as a principled approach to learning biologically consistent developmental trajectories.
@@ G.2 Cell type distributions over time
Figures 57 present spatial maps of transcriptomic datasets across different time points, illustrating how tissue organization and cell-type distributions evolve during development and regeneration. These maps highlight not only changes in cellular composition but also the preservation of spatial neighborhoods and geometrical arrangements of specific cell types over time. Such contextual information, specific to spatial transcriptomics, remains inaccessible to standard flow-matching frameworks. By contrast, ContextFlow is designed to exploit these spatial features, enabling the inference of trajectories that are both temporally smooth and spatially coherent.
@@ G.3 Ligand Receptor Interactions
Based on the activation of NPTX2–NPTXR in Figure 8, we observe that the corresponding communication pattern naturally biases the optimal couplings towards transitions such as Immature dpEX $\rightarrow$ dpEX and Immature nptxEX $\rightarrow$ nptxEX (Figure 9). These transitions are biologically plausible, as they preserve cell type identity within excitatory neuronal lineages while reflecting maturation within the same functional context. This example highlights the richness of the contextual information captured by our proposed biological prior and demonstrates how incorporating such ligand–receptor–driven cues into the coupling process yields more interpretable, biologically consistent trajectories.
Figure 9: Visual translation of the bias that NPTX2–NPTXR LR pattern provides in terms of cell type coupling for two consecutive slides.